3) A researcher documents the number of individuals who have the following genotypes: E1E2 24 E2E2 115 642 Test whether each of the genotypes are in Hardy-Weinberg equilibrium and determine the observed and expected genotype frequencies to fill out the following table: Observed: Expected: Is this gene undergoing evolution? Explain your reasoning. EEEE2EE2
Answer #1

consider allele frequency E1 = p and allele frequency E2 =q

Now total population = 642+24+115 = 781

we can calculate genotype from given data as p = (642+ (24/2))/781 = 0.84 and q= (115+(24/2))/781 = 0.16

here, sum of genotypic frequency is 1 i.e. p+q=1

now phenotypic frequency will be

E1E1 = p2 = 0.842 = 0.7056

E1E2 = 2pq = 2x0.84x0.16= 0.2688

E2E2= q2 = 0.162 = 0.0256

so expected number of population will be

E1E1 = 781x 0.7056 = 551

E1E2 = 781x0.2688 = 210

E2E2 = 781x0.0256=20

from ovserver number of individual and expected number of individual of each phenotype it seems like there is genetic drift in which heterogygous population is getting diminished with time. therefore, there are chances that gene is going under evolution.